# GET: Unlocking the Multi-modal Potential of CLIP for Generalized Category Discovery

### v1

@@ 5.3 Ablation study and analysis
GCD [36] and SimGCD [40] propose to build the classifier on post-backbone features instead of post-projector. Because the ViT backbone of CLIP contains a lot of knowledge learned from substantial image-text pairs, and the projector plays a role in modal alignment, it’s essential to compare the effects of different ViT finetune strategies. As shown in Fig. 4, we conduct multiple evaluations with last-block fine-tuning, projector fine-tuning, and adapter [12] fine-tuning strategies. Though simply fine-tuning the projector can gain a higher accuracy across CUB and Aircraft datasets, it falls behind the last-block fine-tuning method for generic datasets. Overall, our GET perfoms the best among all methods. For a fair comparison, we select the last-backbone fine-tuning strategy for baseline methods and our dual-branch multi-modal learning across all datasets except projector fine-tuning for ImageNet-1K. The detailed results of different fine-tuning strategies are posted in SM.
5.4 Different pre-trained models
In this section, we conduct extensive experiments to investigate how different types of backbones influence the GCD task. As shown in Fig. 5, we evaluate the results of GCD and SimGCD across different types of pretraining models, including DINO [4], which is based on teacher-student learning; MoCo v3 [6], based on contrastive learning; iBOT [47], based on masked image modeling; and CLIP [32], which is based on vision-language contrastive learning. Different types of backbones exhibit varying biases across different datasets, classes, and even paradigms. A key observation is that though promising results have been achieved, different backbones, even powerful CLIP, still perform inferiorly on distinguishing certain visually similar classes, such as the classes in all fine-grained datasets. We argue that this is due to current methods only utilizing a single visual modality of information, another modality may potentially compensate for the lack of discriminative ability. Due to CLIP’s multi-modal training strategy and its strong generalization ability, we decide to introduce it into the GCD task. This not only unleashes the latent potential performance of existing methods but also serves as a bridge for us to leverage multi-modal information. Detailed analysis and additional visualization results are available in SM.
6 Conclusions
In this paper, we propose to leverage multi-modal information to solve the GCD task. In particular, we introduce a text embedding synthesizer to generate pseudo text embeddings for unlabelled data. Our TES module makes it possible to use CLIP’s text encoder, thus unlocking the multi-modal potential of CLIP for the GCD task. Meanwhile, we use a dual-branch training strategy with a cross-modal instance consistency objective, which facilitates collaborative action and mutual learning between the different modalities, thus generating more discriminative classification prototypes. The superior performance on multiple benchmarks demonstrates the effectiveness of our method.
7 Acknowledgments
Supplementary Materials
The Supplementary Materials includes the following sections. First, we introduce our dataset details and training details in Sec. 0.A. Then, we provide more experiment results and analysis in Sec. 0.B. We further discuss the utilization of CLIP in GCD and the intuition behind our TES in Sec. 0.C, and we prove that our TES can deal with a scenario where CLIP lacks information on a specific category class. We provide the pseudo-code in Sec. 0.D. Sec. 0.E shows the visualization of different pre-trained models while Sec. 0.F shows the qualitative results of our method. Finally, we provide the broader impact and limitations in Sec. 0.G.
Appendix 0.A Implementation Details
0.A.1 Dataset details
As shown in Tab. 6, we evaluate our method on multiple benchmarks. We introduced ImageNet-R [15] to the GCD task to demonstrate that GCD based on multi-modal features can generalize across different domains. For ImageNet-R, we subsample the first 100 classes as old classes, leaving the rest as new classes; the labeled dataset $\mathcal{D}_{l}$ consists of half of the old class samples, while the other half and all the new class samples are used to construct unlabelled dataset $\mathcal{D}_{u}$. As for other benchmarks, we follow the previous [36, 40] to sample $\mathcal{D}_{l}$ and $\mathcal{D}_{u}$. ImageNet-R contains various renditions (including art, cartoons, deviantart, graffiti, embroidery, graphics, origami, paintings, patterns, plastic objects, plush objects, sculptures, sketches, tattoos, toys, and video game.) of 200 ImageNet classes resulting in 30,000 images. The statistics for ImageNet-R are shown in Fig. 6. In our experiment, we only sample classes from the set of categories in ImageNet-R to construct labeled dataset $\mathcal{D}_{l}$ and unlabelled dataset $\mathcal{D}_{u}$, without sampling domains. That means, whether old classes or new classes, data from the same category might have multiple domains.
Table 6: The details of the datasets we evaluate on.
Labelled Unlabelled
Dataset Images Classes Images Classes
CIFAR10 [20] 12.5K 5 37.5K 10
CIFAR100 [20] 20.0K 80 30.0K 100
ImageNet-100 [8] 31.9K 50 95.3K 100
CUB [39] 1.5K 100 4.5K 200
Stanford Cars [19] 2.0K 98 6.1K 196
FGVC-Aircraft [27] 1.7K 50 5.0K 100
Herbarium 19 [35] 8.9K 341 25.4K 683
ImageNet-R [15] 7.7K 100 22.3K 200
ImageNet-1K [8] 321K 500 960K 1000
Figure 6: The statistics for ImageNet-R.
0.A.2 Training details
Following [36] and [40], we only fine-tune the last transformer block of the image encoder for all datasets resulting in features with a dimension of 768. The exception is ImageNet-1K, we only fine-tune the last projection layer, resulting in features with a dimension of 512. When fine-tuning the last block of the CLIP visual encoder for ImageNet-1K, there is a risk of gradient explosion. Meanwhile, fine-tuning the last projection layer instead can lead to better results than fine-tuning the last block using a lower learning rate. Additionally, fine-tuning the last projection layer requires fewer computational resources. Therefore, we chose to fine-tune only the last projection layer for ImageNet-1K. The batch size is fixed to 128 for training and 256 for testing. Training is done with an SGD optimizer and an initial learning rate of 0.1 decayed by a cosine annealing rule. We train for 200 epochs on each dataset in both two stages. In the first stage, we set the number of pseudo text tokens to 7. The balance coefficient $\lambda$ is set to 0.35 as [36], while $\lambda_{h}$ and $\lambda_{c}$ are set to 2 and 1, respectively. The temperature value $\tau_{a}$ is set to 0.01 while other temperature values $\tau_{c}$, $\tau_{s}$ and $\tau_{t}$ are as same as [40]. All experiments are conducted with 4 NVIDIA GeForce RTX 3090 GPUs.
Appendix 0.B Additional Experiment Results and Analysis
0.B.1 Error bars for main results
The experimental results presented in the paper are the averages of three independent repeated runs. We provide the performance standard deviation of our main results on all evaluation datasets with three runs in Tab. 7.
Table 7: The standard deviation of our method.
Dataset All Old New
CIFAR10 97.2$\pm$0.1 94.6$\pm$0.1 98.5$\pm$0.1
CIFAR100 82.1$\pm$0.4 85.5$\pm$0.5 75.5$\pm$0.5
ImageNet-100 91.7$\pm$0.3 95.7$\pm$0.0 89.7$\pm$0.4
CUB 77.0$\pm$0.5 78.1$\pm$1.6 76.4$\pm$1.2
Stanford Cars 78.5$\pm$1.3 86.8$\pm$1.5 74.5$\pm$2.2
FGVC-Aircraft 58.9$\pm$1.2 59.6$\pm$0.6 58.5$\pm$1.8
Herbarium 19 49.7$\pm$0.4 64.5$\pm$0.8 41.7$\pm$0.8
ImageNet-1K 62.4$\pm$0.0 74.0$\pm$0.2 56.6$\pm$0.1
ImageNet-R 58.1$\pm$2.4 78.8$\pm$0.5 47.0$\pm$3.9
0.B.2 Additional baseline results
As shown in Tab. 8. We provide results of PromptCAL-CLIP on three fine-grained datasets and three image classification generic datasets. For three fine-grained datasets, our method outperforms PromptCAL-CLIP on all datasets and classes. In particular, we surpass PromptCAL-CLIP by 11.5%, 4.5%, and 4.4% on ‘All’ classes of CUB, Stanford Cars, and Aircraft, respectively. As for the generic datasets, our method surpasses PromptCAL-CLIP on all datasets and achieves the best results on CIFAR-100 and ImageNet-100 datasets.
Table 8: Results of PromptCAL-CLIP.
CUB Stanford Cars FGVC-Aircraft CIFAR10 CIFAR100 ImageNet-100
Methods All Old New All Old New All Old New All Old New All Old New All Old New
PromptCAL-CLIP 65.5 68.7 63.9 74.0 80.8 70.8 54.5 61.8 51.0 88.7 96.5 84.8 80.5 82.4 76.8 87.4 93.6 84.3
0.B.3 Results of two branches
We report the results of visual and text branches for ‘All’ classes across six datasets in Tab. 9. For 2 generic datasets (CIFAR10 and ImageNet-100), though the text branch does not achieve state-of-the-art performance, it still exhibits great performance. For 2 fine-grained datasets (CUB and Stanford Cars), both visual and text branches outperform previous methods by a large margin, while the visual branch performs better. For 2 challenging datasets (ImageNet-1K and ImageNet-R), both visual and text branches achieve remarkable results. Due to the challenging datasets comprising a significant number of unknown classes (ImageNet-1k dataset) or diverse visual concepts within the same class (ImageNet-R dataset), the consistency in text information for the same class contributes to the potentially higher discriminative power of the text branch, leading to a better performance for text branch.
Table 9: The results of two branches.
Dataset Visual Branch Text Branch
CIFAR10 97.2$\pm$0.1 95.1$\pm$0.0
ImageNet-100 91.7$\pm$0.3 90.1$\pm$0.1
CUB 77.0$\pm$0.5 73.6$\pm$0.8
Stanford Cars 78.5$\pm$1.3 73.1$\pm$0.6
ImageNet-1K 62.4$\pm$0.0 63.5$\pm$0.1
ImageNet-R 58.1$\pm$2.4 62.6$\pm$0.9
We also provide the performance evolution of two branches throughout the model learning process on the CUB dataset (see in  Fig. 7), the mutual promotion and fusion of the two branches resulted in excellent outcomes. In our experiments, we consistently and simply select the results from the visual branch.
Figure 7: Performance evolution of two branches throughout the model learning process on the CUB dataset
0.B.4 Estimation of class number in unlabelled data.
Vaze et al. [36] provides an off-the-shelf method to estimate the class number in unlabelled data. We introduce text embeddings generated by our TES into the off-the-shelf method by simply concatenating text and image features before class number Estimation. As shown in Tab. 10, multi-modal features can estimate a more accurate class number than single visual features, demonstrating our multi-modal method is effective in category number estimation.
Table 10: Estimation of class number in unlabelled data. We use an off-the-shelf method provided by Vaze et al. [36] to estimate the class number in unlabelled data. The table shows the estimated number and the error.
Method CIFAR-10 CIFAR-100 ImageNet-100 CUB-200 SCars
Ground truth 10 100 100 200 196
GCD-CLIP 5 (50%) 94 (6%) 116 (16%) 212 (12%) 234 (19%)
$+$TES 8 (20%) 97 (3%) 109 (9%) 212 (12%) 220 (12%)
0.B.5 Results of different ViT fine-tuning strategies
We reported the detailed results of different fine-tuning strategies based on SimGCD-CLIP in Tab. 11. Only fine-tuning an additional adapter proved insufficient to fully leverage the performance of the pre-trained model, hence yielding the poorest results. While fine-tuning the projection layer can achieve good results with relatively low computational resources on the CUB and Stanford Cars datasets, its performance on other datasets is significantly inferior to fine-tuning the last block of the model, thus we choose the last-block fine-tuning strategy for baseline methods on most datasets.
Table 11: Results of different ViT fine-tuning strategies.
CUB Stanford Cars FGVC-Aircraft CIFAR10 CIFAR100 ImageNet-100
Methods All Old New All Old New All Old New All Old New All Old New All Old New
last-block 71.7 76.5 69.4 70.0 83.4 63.5 54.3 58.4 52.2 97.0 94.2 98.4 81.1 85.0 73.3 90.8 95.5 88.5
projector 72.2 76.3 70.1 71.5 83.0 65.9 44.6 52.6 40.5 95.6 96.3 95.2 77.1 82.3 66.7 88.7 93.9 86.0
adapter 63.9 72.5 59.5 62.1 72.9 56.9 37.1 40.7 35.3 95.6 96.4 95.2 74.2 80.6 61.6 84.3 93.2 79.8
0.B.6 Computation complexity analysis.
Tab. 12 shows the computation complexity. Our TES uses a frozen visual encoder and stage 2 finetunes the last block in another visual encoder, thus the 2 stages share the same visual encoder for the first 11 blocks but a different last block, resulting in a low computational complexity increase.
Table 12: Computation complexity analysis
Inference Time Learnable Params FLOPs
Methods (s/per img) (M) (G)
SimGCD-CLIP $\text{5.2}\times\text{10}^{-\text{3}}$ 13.4 35.2
GET(ours) $\text{5.2}\times\text{10}^{-\text{3}}$ 15.6 38.6
0.B.7 Analysis of loss components in TES
The objective of our TES module contains an align loss on all samples and a distill loss on labeled samples. The align loss $\mathcal{L}_{align}$ leverages the modality alignment property of CLIP’s encoders, aligning the generated pseudo-text embeddings with their corresponding visual feature. The distill loss $\mathcal{L}_{distill}$ is used to guide pseudo text embeddings of the network’s output towards the real semantic corresponding space and adapt the model to the distribution of the dataset. As shown in Tab. 13, utilizing either distillation loss or alignment loss independently in TES followed by a dual-branch training does not yield performance as strong as the combination of both losses.
Table 13: Ablation study of loss components in TES.
$\mathcal{L}_{align}$ $\mathcal{L}_{distill}$ CUB
All Old New
(1) 74.7 76.9 73.5
(2) 75.3 77.5 74.2
(3) 77.0 78.1 76.4
0.B.8 Experiments on pseudo-text token quantities in TES
We set the number of pseudo-text tokens in TES to 7 across all datasets. Experiment results on token quantities for the CUB dataset are presented in Fig. 8.
Figure 8: Experiment results on token quantities for the CUB dataset.
0.B.9 Experiments on the Clevr-4 dataset
Recently, [38] presented a synthetic dataset Clevr-4 to examine whether the GCD method can extrapolate the taxonomy specified by the labeled set. Most attributes of Clevr-4, such as shape, color, and count, are easily clustered (achieving close to 99% accuracy with CLIP). However, texture attributes pose a certain level of challenge. Therefore, we evaluate our method on the texture attributes of Clevr-4. As shown in Tab. 14, our method achieves higher accuracy and lower standard deviation compared to SimGCD-CLIP, proving that the GCD method can cluster data at specified levels based on the constraint of labeled text information, which is worthy of attention and exploration.
Table 14: The results on Clevr-4 (Texture) in 5 runs.
Methods All Old New
SimGCD-CLIP 83.1$\pm$7.4 99.2$\pm$0.3 75.1$\pm$10.9
GET(ours) 90.0$\pm$1.9 99.2$\pm$0.2 85.5$\pm$2.8
Appendix 0.C Discussion about CLIP in GCD and our TES
@@
Furthermore, we try to prove our TES can deal with a scenario where CLIP lacks information on a specific category class. Since CLIP saw most of the visual concepts and corresponding texts before 2022, we constructed a small dataset of new energy vehicles (NEV) that appeared in 2023. The NEV dataset contains 12 categories, each with 50 images from the Internet, and the classnames (Tab. 15) of the dataset consist of the brand and model of the car. We split them in the same way as standard benchmarks.
Table 15: The classnames for the NEV dataset.
@@
As shown in Tab. 16, due to the lack of information on the NEV dataset in CLIP, achieving only 10% zero-shot accuracy across all classes and CLIP can only recognize 5 out of the 12 categories. Due to the dataset’s fine-grained nature and few samples, SimGCD with the DINO backbone performs poorly on ‘New’ categories. Although fine-tuning with CLIP’s visual encoder as the backbone can yield great results, our GET with multi-modal features demonstrates a 9.7% improvement on ‘New’ categories. This proves the effectiveness of our TES in scenarios where CLIP lacks information or the CLIP text encoder has not encountered corresponding class names.
Table 16: The results on the NEV dataset.
@@
The intuition behind our TES can be explained from two perspectives, as illustrated in Fig. 9. The first perspective is that our trained TES can be considered as a special fine-tuned text encoder. This text encoder takes visual images as input and produces corresponding textual features as output. Our align loss ensures modal alignment, while the distill loss facilitates the model’s adaptation to the dataset’s distribution. The second perspective is that TES can be regarded as a caption model [28]. For each input image, TES assigns a corresponding caption, expressing each caption in the form of text features. The text embeddings or captions corresponding to images can serve as valuable supplementary information, assisting the GCD task in a multimodal manner.
Figure 9: The intuition behind our TES. (a) Well-trained TES module; (b) TES can be considered as a special fine-tuned text encoder; (c) TES can be regarded as a caption model.
Appendix 0.D Pseudo-code
We provide a pseudo-code of our method in Algorithm 1. Since information from different modalities is exchanged and learned through CICO and injected into the trainable visual backbone, we utilize the last-epoch visual branch for inference. Specifically, passing unlabeled data through the trained visual encoder $f_{v}(\cdot|\theta_{\text{v}})$ and classifier $\eta(\cdot|\theta_{\text{c}})$ followed by the Hungarian algorithm can obtain the clustering results.
Input: Training dataset $\mathcal{D}$ contains labeled dataset $\mathcal{D}_{l}$ and unlabelled dataset $\mathcal{D}_{u}$, a FC layer $l(\cdot|\theta_{\text{t}})$ and a MLP layer $g(\cdot|\theta_{\text{m}})$, CLIP’s image encoder $f_{v}(\cdot|\theta_{\text{v}})$ and text encoder $f_{t}(\cdot)$, a classifier $\eta(\cdot|\theta_{\text{c}})$ and a linear projection $p(\cdot|\theta_{\text{p}})$.
Output: Trained backbone $f_{v}(\cdot|\theta_{\text{v}})$ and classifier $\eta(\cdot|\theta_{\text{c}})$.
/* Stage 1: TES Training */
1 for each epoch $e$=1…$E_{1}$  do
2       for  $(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\in\text{each batch}$ do
             $\boldsymbol{z}_{i}^{v}=f_{v}(\boldsymbol{x}_{i})$ // visual embedding
3
            $\boldsymbol{t}_{i}=l(\boldsymbol{z}_{i}^{v}))$ // pseudo text tokens
4
            $\boldsymbol{\hat{z}}_{i}^{t}=f_{t}(\boldsymbol{t}_{i})$ // pseudo text embedding
5
6            Compute $\mathcal{L}_{align}$ by Eqs. (5,6) on $\boldsymbol{z}_{i}^{v},\ \boldsymbol{\hat{z}}_{i}^{t}$
7            Compute $\mathcal{L}_{distill}$ by Eq. (7) for labeled data
8            Compute overall TES loss $\mathcal{L}_{TES}$ by Eq. (8)
9            Back-propagation and optimize $\theta_{\text{t}}$.
10
/* Stage 2: Dual-branch training */
11 Trained $TES(\cdot)$
12for each epoch $e$=1…$E_{2}$ do
13       for  $(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\in\text{each batch}$ do
             /* Visual-branch */
14             $\boldsymbol{z}_{i}^{v}=f_{v}(\boldsymbol{x}_{i}),\ \boldsymbol{h}_{i}^{v}=g(\boldsymbol{z}_{i}^{v}),\ \boldsymbol{p}_{i}^{v}=\eta(\boldsymbol{z}_{i}^{v})$
15            Replace $\boldsymbol{h}$ in Eqs. (1,2) with $\boldsymbol{h}^{v}$ for $\mathcal{L}_{ucon}^{v}$ and $\mathcal{L}_{scon}^{v}$
16            Compute $\mathcal{L}_{\text{rep}}^{v}$ by Eq. (9)
17            Replace $\boldsymbol{p}$ in Eqs. (3,4) with $\boldsymbol{p}^{v}$ and compute $\mathcal{L}_{\text{cls}}^{v}$
18            Compute $\mathcal{L}_{\text{dual}}^{v}$ by Eqs. (10)
            /* text-branch */
             $\boldsymbol{\hat{z}}_{i}^{t}=TES(\boldsymbol{x}_{i})$ // pseudo text embedding
19
20            $\boldsymbol{\hat{z}}_{i}^{tl}=p(\boldsymbol{\hat{z}}_{i}^{t}),\ \boldsymbol{h}_{i}^{t}=g(\boldsymbol{\hat{z}}_{i}^{tl}),\ \boldsymbol{p}_{i}^{t}=\eta(\boldsymbol{\hat{z}}_{i}^{tl})$
21            Compute $\mathcal{L}_{\text{dual}}^{t}$ by Eq. (11)
22            Compute the multi-modal mean entropy regularization $H_{mm}$
            /* CICO */
23             Calculate the visual and text anchors $\mathcal{P}_{v},\mathcal{P}ICO_{t}$
24            Compute the instance relationships by Eqs. (12,13)
25            Compute $\mathcal{L}_{\text{CICO}}$ by Eq. (14) and the overall loss $\mathcal{L}_{\text{Dual}}$ by Eq. (15)
26            Back-propagation and optimize $\theta_{\text{v}},\ \theta_{\text{m}},\ \theta_{\text{c}},\ \theta_{\text{p}}$.
27
28return $f_{v}(\cdot|\theta_{\text{v}})$, $\eta(\cdot|\theta_{\text{c}})$
Algorithm 1 Pseudocode for  GET.
Appendix 0.E Visualization of Different Pre-trained Models
In this paper, we perform an extensive empirical investigation to explore the impact of different types of pre-trained models on GCD clustering performance, which clearly demonstrates that different types of backbones exhibit varying biases across different datasets, classes, and even paradigms. We visualize and compare the attention map of class tokens of different backbones in  Fig. 10. For the CUB dataset, the DINO, iBOT, and MoCO backbones tend to focus more on the feathers of the birds, while CLIP additionally emphasizes the more discriminative head area. For the StanfordCars dataset, the DINO backbone focuses on the car light and wheel; the MoCo backbone focuses on the front fenders of the car, which is less discriminative; the iBOT backbone focuses on the car light and the car window, which is more discriminative than DINO thus leading to better results; the CLIP backbone focus on both the front of the car and global information, showcasing stronger discriminative capabilities.

### v2

@@ 5.3 Ablation study and analysis
Different ViT fine-tuning strategies.  GCD [3] and SimGCD [4] propose to build the classifier on post-backbone features instead of post-projector. Because the ViT backbone of CLIP contains a lot of knowledge learned from substantial image-text pairs, and the projector plays a role in modal alignment, it’s essential to compare the effects of different ViT finetune strategies. As shown in Fig. 3, we conduct multiple evaluations with last-block fine-tuning, projector fine-tuning, and adapter [44] fine-tuning strategies. Though simply fine-tuning the projector can gain a higher accuracy across CUB and Aircraft datasets, it falls behind the last-block fine-tuning method for generic datasets. Overall, our GET perfoms the best among all methods. For a fair comparison, we select the last-backbone fine-tuning strategy for baseline methods and our dual-branch multi-modal learning across all datasets except projector fine-tuning for ImageNet-1K.
Effectiveness of text embedding synthesizer.  In order to prove that our text embedding synthesizer can generate reliable and discriminative representations, we visualize the text embeddings of CIFAR10 with t-SNE. As shown in Fig. 4, the initial text embeddings within the same class exhibit clear clustering, and the learnable embeddings further produce compacter clusters. Moreover, we introduce TES into the non-parametric GCD by straightforwardly concatenating text and image features before semi-supervised k-means classification. As in Tab. VI, with the help of text information, GCD gains about 5% average improvement on ‘All’ classes over 3 datasets, demonstrating the importance of multi-modal information in GCD task and our TES can be widely used in multiple GCD methods.
TABLE VI: Effectiveness of TES in non-parametric GCD.
FGVC-Aircraft ImageNet-100 ImageNet-R
Method All Old New All Old New All Old New
GCD-CLIP 45.3 44.4 45.8 75.8 87.3 70.0 44.3 79.0 25.8
+TES 49.6 49.3 49.8 80.0 95.1 72.4 49.4 79.4 33.5
In TES, we use a single linear layer to transform the visual embedding to pseudo tokens that serve as input to the CLIP’s text encoder. We set the number of pseudo text tokens to 7 across all datasets. Experiment results on the number of tokens and layers in TES for the CUB dataset are presented in Fig. 5, proving that a single linear layer can effectively transfer visual features into text tokens while reducing the computational cost.
Figure 5: Experiment results on the number of tokens and layers in TES for the CUB dataset.
TABLE VII: Experiments on different methods for generating pseudo text embeddings.
Method All Old New
Baseline SimGCD-CLIP 71.7 76.5 69.4
Text-Retrieval WordNet 69.8 77.1 66.2
CC3M 72.3 79.1 68.9
Text-Generation BLIP 67.1 74.3 63.5
BLIP-2 71.3 73.5 70.2
Feature-Generation TES (Ours) 77.0 78.1 76.4
TES w/o $\mathcal{L}_{align}$ 74.7 76.9 73.5
TES w/o $\mathcal{L}_{distill}$ 75.3 77.5 74.2
Our TES generates pseudo text embeddings for unlabelled data in a feature-generation approach. To demonstrate the superiority of TES, we conduct experiments that replaced the embeddings generated by TES with embeddings obtained through text-retrieval or text-generation on the CUB dataset. For text-retrieval, we retrieve the most similar text for each image in the dataset from the corpus (WordNet [45] with 117k names and CC3M [46] with 3M captions) based on the cosine similarity of the image and text embeddings. For text-generation, we use BLIP [47] and BLIP-2 [48] to perform the VQA task using the question “What’s the name of the bird in the image?”, generating text answer for each image. After obtaining the corresponding text, we encode it into text embeddings and conduct dual-branch training. As in Table VII, for fine-grained datasets like CUB, text-retrieval and text-generation methods do not enhance performance and can even be harmful. This is due to the high visual similarity in fine-grained images, leading to lower accuracy in the obtained texts. As a result, the generated text embeddings are less discriminative. Our TES, from the perspective of feature generation, optimizes through effective $\mathcal{L}_{align}$ and $\mathcal{L}_{distill}$, generating discriminative text features for the images.
5.4 Discussion about using CLIP in GCD
@@
TABLE VIII: The classnames for the NEV dataset.
@@
Furthermore, we conduct a toy experiment to prove our TES can deal with a scenario where CLIP lacks information on a specific category class. Since CLIP saw most of the visual concepts and corresponding texts before 2022, we constructed a small dataset of new energy vehicles (NEV) that appeared in 2023. As shown in Table VIII, the NEV dataset contains 12 categories, each with 50 images from the Internet, and the classnames of the dataset consist of the brand and model of the car. We split them in the same way as standard benchmarks.
TABLE IX: The results (%) on the NEV dataset.
@@
As shown in Table IX, due to the lack of information on the NEV dataset in CLIP, achieving only 10% zero-shot accuracy across all classes and CLIP can only recognize 5 out of the 12 categories. Due to the dataset’s fine-grained nature and few samples, SimGCD with the DINO backbone performs poorly on ‘New’ categories. Although fine-tuning with CLIP’s visual encoder as the backbone can yield great results, our GET with multi-modal features demonstrates a 9.7% improvement on ‘New’ categories. This proves the effectiveness of our TES in scenarios where CLIP lacks information or the CLIP text encoder has not encountered corresponding class names.
@@
Our trained TES can be considered as a special fine-tuned text encoder. This text encoder takes visual images as input and produces corresponding textual features as output. Our align loss ensures modal alignment, while the distill loss facilitates the model’s adaptation to the dataset’s distribution. For classes that CLIP has not seen before, TES can be regarded as a caption model [10]. For each input image, TES assigns a corresponding caption, expressing each caption in the form of modal-aligned text features. The text embeddings or captions corresponding to images can serve as valuable supplementary information, assisting the GCD task in a multimodal manner.
5.5 Different pre-trained models
In this section, we perform an extensive empirical investigation to explore the impact of different types of pre-trained models on GCD clustering performance, which clearly demonstrates that different types of backbones exhibit varying biases across different datasets, classes, and even paradigms. We choose DINO [8], which is based on teacher-student learning; MoCo v3 [49], based on contrastive learning; iBOT [50], based on contrastive masked image modeling; and CLIP [9], which is based on vision-language contrastive learning.
We first evaluate the results of GCD and SimGCD across different types of pretraining models. As shown in Fig. 6, different types of backbones exhibit varying biases across different datasets, classes, and even paradigms. For example, iBOT outperforms DINO in non-parametric GCD, but DINO excels in parametric GCD. MOCO demonstrates the strongest category discovery ability for the CIFAR dataset. CLIP performs exceptionally well across all datasets, yet struggles with low-resolution CIFAR data in parametric GCD.
We then visualize and compare the attention map of class tokens of different backbones in  Fig. 7. For the CUB dataset, the DINO, iBOT, and MoCO backbones tend to focus more on the feathers of the birds, while CLIP additionally emphasizes the more discriminative head area. For the StanfordCars dataset, the DINO backbone focuses on the car light and wheel; the MoCo backbone focuses on the front fenders of the car, which is less discriminative; the iBOT backbone focuses on the car light and the car window, which is more discriminative than DINO thus leading to better results; the CLIP backbone focus on both the front of the car and global information, showcasing stronger discriminative capabilities.